# Generative Photomontage

### v2

@@ 6 Discussion and Limitations
In this work, we proposed a new approach for generating images: by compositing it from multiple ControlNet-generated images. To this end, we introduced a novel method for segmenting and blending the images using their diffusion features, and we showed many compelling applications. At a broader level, our work suggests a new user workflow for interacting with text-to-image models: rather than trying to get the model to output the final end-product (i.e., a single image that contains everything the user wants), we treat the model’s output as intermediate outputs, from which users provide further input to create their final end-product. This approach not only gives users more fine-grained control over the final output, but also allows us to fully utilize the model’s generative capabilities in creating diverse candidates. We hope this work inspires new ways of interacting with generative models.
Limitations.   While we have shown our method’s versatility in various applications, we also observe several limitations. First, our current graph cut parameters are empirically chosen to encourage congruous regions, which penalizes seam circumference. While this works well for many cases, if the target object has a curvy outline, it may require additional user strokes to obtain a finer boundary (Figure 12). Since graph cut is solved in near real-time (${\sim}1$s), users can quickly check the graph-cut result by visualizing it in image space and iterate as needed.
Second, our method assumes some spatial consistency among images in the stack. If the images differ significantly in scene structure, it will rely more on the user to select proper regions to form a valid scene (Figure 13a). Alternatively, users can increase spatial consistency by adding more structure to the input control (Figure 13b). Future work can investigate ways to relax spatial consistency constraints and automatically account for dramatic scene structure changes during segmentation and blending.
@@
Acknowledgments   We are grateful to Kangle Deng for his help with setting up the user survey. We also thank Maxwell Jones, Gaurav Parmar, and Sheng-Yu Wang for helpful comments and suggestions and Or Patashnik for initial discussions. This project is partly supported by the Amazon Faculty Research Award, DARPA ECOLE, the Packard Fellowship, and a joint NSFC-ISF Research Grant no. 3077/23.
Appendix A Additional Results.
We include additional results in Figure 15 and 16. Figure 15 is an additional example of appearance mixing, where the user combines a green bird with blue wings and a yellow beak. Figure 16 is an additional example of improving prompt alignment. As shown, vanilla ControlNet struggles to correctly adhere to the long, complicated prompt. Using our method, the user can break the prompt up into multiple, short prompts and composite the images generated from the short prompts to form their desired image.
Appendix B Additional Baseline Comparisons
@@
Here, we qualitatively compare our method with additional baselines, shown and summarized in Figure 17 and 18. Our method outperforms the baselines in terms of preserving local appearances and blending harmoniously. Below, we describe implementation details.
CollageDiffusion [64].   We followed their released demo for generating baseline results. Their method allows users to tweak parameters per image layer and scene, such as added noise level and cross attention modulation, and then outputs a composite. We created the input image layers by resizing the masks from feature-space graph-cut to image space and applying them to each image. Reducing the added noise to zero preserves all local regions but with little harmonization (similar to copy-paste). To minimize changes, we used a set of parameters that preserves the background region (base image) but allows the foreground injected regions to change. We set noise levels to $0.05$ for the base image and $0.4$ for the non-base images. We set noise blur to $30$ to smooth the seams. To ensure spatial fidelity of the image regions, we set the cross attention modulation of non-base images to $0.5$, following their released examples. We also use their textual inversion feature and learn a special embedding for each image layer. If there are multiple image layers (regions) that correspond to the same word (e.g., multiple robot parts correspond to “robot” in the prompt “A robot from the future”), we choose the one that most closely matches the word description.
Deep Image Blending [83].   We ran their released code to generate the baseline results. Their method takes in two images to composite (a source and target image) as well as a mask of the source image. To create the mask for each image, we resized the corresponding graph-cut mask from feature to image space. For composites of more than two input images, we iteratively ran Deep Image Blending on pairs of images to build towards the final composite. One input constraint is that their method takes square images as input, so we pre-process our image stack by resizing them to square images and resize the output back to the original aspect ratios.
@@
Blended Latent Diffusion + MultiDiffusion [4, 8].   We experimented with a modified version of Blended Latent Diffusion, where we fused the noises of different image regions with greater overlap, inspired by MultiDiffusion. Specifically, we dilate the mask of each region (in feature space) with a $3\times 3$ kernel and average the noise within the overlapped regions. As shown in Figure 17, this leads to artifacts near the seams (site of overlap) or changes local appearances.
@@
GP-GAN [79].   We ran their released code with their pre-trained model to generate the results. GP-GAN takes in two images to composite (a source and destination image), along with a mask of the source image to insert into the destination image. For composites of more than two images, we iteratively ran GP-GAN on pairs of images to build towards the final composite. We created the input masks by resizing the diffusion-space graph-cut results to the original resolution of each image.
@@
Appendix C Additional Ablations
Here, we show additional ablation results on our graph cut and feature injection.
C.1 Graph Cut
Our method uses the $K$ features from self-attention layers to compute the graph cut energy terms, i.e., the pairwise terms which dictate seam costs. Here, we consider alternative features (Figure 19).
$Q$ and $V$ features.   $Q$ features lead to similar segmentations as $K$ features but may exhibit over-and under-segmentations near boundaries (circled in Figure 19). $V$ features generally do not lead to seams that align with semantic boundaries.
DINOv2 Features.   We also experimented with DINOv2 features [54] in place of diffusion features. Specifically, we used the small DINOv2 model, which outputs features per $14\times 14$ patch. For fair comparison, we resized all images such that the size of output features would equal that of $K$ features. As shown in Figure 19, if we directly apply our graph cut method (which uses the features’ top-10 PCA components), we see errors in segmentation. If we use the top-100 PCA components, we get better segmentations for most examples but not all (e.g., the dog in Figure 19d).
Discussion.   While we chose to go with $K$ features due to the best segmentation, we found that our feature blending method is robust to small errors in segmentation boundaries. For example, the segmentations we get from $Q$ features are similar to that of $K$ features. While $Q$ features may result in over- or under-segmentations near boundaries, our feature blending method blends the seams well, such that the difference in the resulting images is not too noticeable (Figure 20).
C.2 Feature Injection
Our method injects the composite $Q^{\text{comp}}$, $K^{\text{comp}}$, $V^{\text{comp}}$ into the U-Net’s self-attention maps for all layers and time steps. Following related works in self-attention injection [14, 2], we consider injecting only in decoder layers and in later time steps.
Time Steps.   We used a total of $20$ denoising time steps for all our results. Prior work shows that earlier denoising time steps form image layout and shape, while later time steps form detailed appearance [86, 14]. We experimented with applying our injection after $5$, $10$, and $15$ time steps. As shown in Figure 22, starting the injection later tends to create artifacts in the blended output. In most cases, starting the injection after $5$ time steps lead to comparable results to our method (which injects all time steps). However, in scenes that require the image structure to adapt near the seams, doing so may prevent the model from adapting sufficiently. For example, if we inject after $5$ time steps, the shadow from the removed rock at the apple bite still shows up in the final image (circled in red), whereas it is completely removed if we apply the injection for all time steps.
Discussion.   If the required adaption near seams is minor, and the user does not mind changes to local appearances, one can consider applying the injection only in decoder layers (D1-D3) and after $5$ time steps. This will require less storage space for the initially generated $Q,K,V$ feature vectors, as only the features in the relevant layers and time steps will need to be stored. However, if the user wishes to maximize adaptation near seams and to maximize local appearance fidelity, we recommend using our full injection approach.
Appendix D User Survey Details
@@
Figure 14: User Survey (Blending). We show users the input images and selected regions to blend, and then ask them pick the best result from four images: one generated by our method, and the others generated by Digital Photomontage, Blended Latent Diffusion, and MasaCtrl+ControlNet.
@@
We launched two anonymous user surveys among computer science graduate students, in order to compare our method with three other baselines (Interactive Digital Photomontage [1], Blended Latent Diffusion [4], MasaCtrl+ControlNet [14]). For each of the $12$ scenes (Figures 5, 6, 7, 15), we created results using our method and the three baselines. For each scene, we showed the four results side-by-side and asked users to select the best result. The order of the four images are randomized for each scene, and the order of scenes are also randomized in each survey.
@@
In the first survey, we asked “Which of the following images appear most realistic to you?” In the second survey, we showed the input images and asked users to pick the image that is the best at blending them. Specifically, we showed the full input images as well as the graph-cut regions (to be blended) on each input image. To help users focus on the individual image regions, we darken out other parts of the input image (Figure 14). Then, we asked “Which of the following images are BEST at blending all the selected image regions shown on top?”
@@
Users are asked to complete the first survey (realism) before the second survey (blending) to avoid bias. We received $16$ completed surveys for the realism one (total of $192$ responses across all scenes) and $10$ completed surveys for the blending one (total of $120$ responses across all scenes).
Appendix E Result Details
@@
We used various Stable Diffusion pre-trained models (v1.5) to generate results in this paper. Specifically, we used the canny edge model (Figures 5d, 6a,d,e, 7, 15), the scribble model (Figures 5b,c,e, 6b), the depth map model (Figure 5a), the Openpose model (Figure 6c), and the HED model (Figures 13, 16).
@@
The input conditions (e.g., edge map, sketches) are manually created or derived from images released by ControlNet [84] or found on the web [46, 41, 78, 69, 19, 3].
Figure 15: Additional Result: Bird. Users can combine feather colors of birds. (a) Input to vanilla ControlNet. (b) The user selects the green feathers in the first image, the yellow beak and foothold of the second image, and the blue wings from the third image. (c) Generative Photomontage’s graph cut (top) and final output (bottom).

### v3

@@ 6 Discussion
In this work, we proposed a new approach for generating images: by compositing it from multiple ControlNet-generated images. At a broader level, our work suggests a new user workflow for interacting with text-to-image models: rather than trying to get the model to output the final end-product (i.e., a single image that contains everything the user wants, which is often difficult), we treat the model’s output as an intermediate step, from which users provide further input to create their final end-product. Our approach not only gives users more fine-grained control over the final output, but also allows us to fully utilize the model’s generative capabilities in creating diverse candidates. We hope this work inspires new ways of interacting with generative models.
@@
Acknowledgments   We are grateful to Kangle Deng for his help with setting up the user survey. We also thank Maxwell Jones, Gaurav Parmar, Sheng-Yu Wang, and Or Patashnik for helpful comments and suggestions. This project is partly supported by the Amazon Faculty Research Award, DARPA ECOLE, the Packard Fellowship, the IITP grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project), and a joint NSFC-ISF Research Grant no. 3077/23.
Appendix
In Appendix A, we discuss runtime and storage performance of our method. Then, we show more results (Appendix C), qualitative comparisons with baselines (Appendix D), and additional ablations (Appendix E). Next, we provide implementation details of our baselines (Appendix F), the user survey details (Appendix G), and result details (Appendix H). Finally, we discuss our work’s limitations in Appendix I.
Appendix A Runtime and Storage
Our graph cut runs in ${\sim}1$ sec in total, which includes preprocessing (e.g., PCA space computation, computing energy costs) and the actual graph-cut optimization. In theory, the graph-cut optimization depends on the number of variables (feature map resolution) and candidate labels (number of images). In practice, however, there is very little overhead. For our results ($\sim$64x64 feature maps with 2-5 images), the solver takes $\sim$3ms on average to find the solution.
The blending stage takes ${\sim}3$ seconds on an NVIDIA A6000 GPU. For reference, one forward pass of vanilla ControlNet takes ${\sim}2$ seconds on the same GPU. Storage space for the initially generated $QKV$ features depends on image resolution and is ${\sim}2$GB for a $512\times 512$ image.
While we resorted to storing the features on disk, most modern GPUs have enough VRAM to generate multiple images in a single batch, obviating the need to store features on disk. For example, A6000 (48GB) can generate up to $60$ images for SD1.5 ControlNet. Hence, one can re-generate the image stack and its QKV features in the same batch as the composite image during the second pass. Alternatively, one could store a subset of features with a trade-off of lower appearance fidelity (please see Appendix E.3 for more details).
Appendix B Graph Cut Parameters
We use the same graph cut parameters, C = $1010^{6},\lambda=100,\sigma=$ 10, across all our results. $\sigma$ controls how quickly the pairwise cost decreases with feature changes; a large $\sigma$ prevents cuts from aligning with semantic boundaries, and a small $\sigma$ is overly sensitive to feature differences, causing cuts to fall directly on the strokes (Figure 10). $\lambda$ scales the pairwise costs to reduce rounding errors for the solver, and $C=10^{6}$ ensures unary costs dominate, enforcing a hard constraint for strokes.
Appendix C Additional Results
We include our full results in Figures 13-19. We show applications of appearance mixing (Figures 13 and 14), shape and artifacts correction (Figure 15), and prompt alignment (Figures 16, 17, 18 and 19).
Appendix D Additional Baseline Qualitative Comparisons
@@
We show additional qualitative comparisons of our method with baselines in Figures 20, 21, and 22. Our method outperforms the baselines in terms of preserving local appearances and blending harmoniously.
Appendix E Additional Ablations
Here, we show additional ablation results on our graph-cut segmentation, the features used for the graph cut, and the self-attention feature injection.
E.1 Multi-Image Segmentation
In Table 3, we show quantitative results of ablating our graph-cut segmentation with SAM [44] and IDP’s pixel-space graph cut [1]. In both cases, the masked LPIPS and PSNR scores became worse, reflecting a lower fidelity of local appearances due to less accurate segmentations. Please see Section 5.2 and Fig. 3 and 9 for qualitative examples.
Figure 10: Graph Cut: $\sigma$ parameter. A large $\sigma$ prevents cuts from aligning with semantic boundaries, and a small $\sigma$ is overly sensitive to feature differences, causing cuts to fall directly on the strokes. Our graph cut parameter ($\sigma=10$) strikes a balance and segments image regions along semantic boundaries.
Masked LPIPS $\downarrow$ Masked SSIM $\uparrow$ PSNR $\uparrow$ Seam Gradient Score min, avg, max 0.255, 0.340, 0.425
Ours 0.123 0.815 22.46 0.339
w/ SAM 0.223 0.697 17.71 0.325
w/ IDP graph cut 0.146 0.786 21.28 0.312
Table 3: Ablation: Segmentation.
E.2 Graph Cut
Our method uses the output features of the key projection matrix from self-attention layers ($K$ features) to compute the graph cut energy terms, i.e., the pairwise terms that dictate seam costs. Here, we consider alternative features, as discussed in Section 5.2, and show its results in Figure 23. $K$ features performs the best, with $Q$ features a close second. Using $K$ features at the last timestep is also better than averaging $K$ features across earlier timesteps.
$Q$ and $V$ features.   Here, we use output features from query and value projection matrices, namely $Q$ and $V$ features, for segmentation. $Q$ features usually lead to similar segmentations as $K$ features but may exhibit over-and under-segmentations near boundaries in some cases (circled in Figure 23). $V$ features generally do not lead to seams that align with semantic boundaries.
DINOv2 Features.   We also experimented with DINOv2 features [58] in place of diffusion features. Specifically, we used the small DINOv2 model, which outputs features per $14\times 14$ patch. For a fair comparison, we resized all images such that the size of output features would equal that of $K$ features. As shown in Figure 23, if we directly apply our graph cut method (which uses the features’ top-10 PCA components), we see errors in segmentation. If we use the top-100 PCA components, we get better segmentations for most examples but not all (e.g., the dog in Figure 23d).
Discussion.   While we chose to go with $K$ features due to the best segmentation, we found that our feature blending method is robust to small errors in segmentation boundaries. For example, the segmentations we get from $Q$ features may result in over- or under-segmentations near boundaries, but our feature blending method blends the seams well, such that the difference in the resulting images is not too noticeable (Figure 24).
E.3 Feature Injection
Figure 11: Ablation with Alternative Self-Attention Injection Strategies. (a) Input graph-cut segmentation, visualized here by resizing the diffusion-feature masks to match image dimensions and compositing in pixel space. (b) Inspired by StyleAligned [32], we replace $K^{\text{comp}},V^{\text{comp}}$ with $K^{\text{concat}}=[K^{1},K^{2},...,K^{N}]$ and $V^{\text{concat}}=[V^{1},V^{2},...,V^{N}]$. (c) Similar to Equation 4, we injected $K^{\text{model}}$ and $V^{\text{model}}$ for the base image. (d) Our method can preserve local image appearances while harmoniously blending them. Please zoom in for details.
In Sec 5.1, we showed quantitative results of (1) using shared (concatenated) key and value features, i.e, $K^{\text{concat}}=[K^{1},K^{2},...,K^{N}]$ and $V^{\text{concat}}=[V^{1},V^{2},...,V^{N}]$, inspired by Hertz et al. [32]; and (2) using $K^{\text{model}}$ and $V^{\text{model}}$ for the base image, similar to Equation 4. Here, Figure 11 shows qualitative samples of final outputs along with the intermediate feature-based graph cut, visualized in image space.
As shown, the alternative injection schemes can change the local image appearance, such as the dog’s background, the robot’s head, the snake color, and the waffle. Our results align with the observation [2, 15] that $Q$ features influence the image structure, and $K$ and $V$ features influence the image appearance.
$K^{\text{concat}},V^{\text{concat}}$ allows $Q$ features to match with keys that are not from the target image region and can thus result in appearance changes. Our method only makes the target image region’s keys and values available and, hence, is able to preserve local appearances. Using $K^{\text{model}},V^{\text{model}}$ for the base image also allows appearances to change. By injecting $K^{B}$ and $V^{B}$ of the base image stored during initial generation, our method can preserve their appearance during blending.
Our method injects the composite $Q^{\text{comp}}$, $K^{\text{comp}}$, $V^{\text{comp}}$ into the U-Net’s self-attention maps for all layers and time steps. Following related works in self-attention injection [15, 2], we also consider injecting only in decoder layers and in later time steps.
Time Steps.   We used a total of $20$ denoising time steps for all our results. Prior work shows that earlier denoising time steps form image layout and shape, while later time steps form detailed appearance [92, 15]. We experimented with applying our injection after $5$, $10$, and $15$ time steps. As shown in Figure 26, starting the injection later tends to create artifacts in the blended output. In most cases, starting the injection after $5$ time steps leads to comparable results to our method (which injects all time steps). However, in scenes that require the image structure to adapt near the seams, doing so may prevent the model from adapting sufficiently. For example, if we inject after $5$ time steps, the shadow from the removed rock at the apple bite still shows up in the final image (circled in red), whereas it is completely removed if we apply the injection for all time steps.
Discussion.   If the required adaption near seams is minor, and the user does not mind changes to local appearances, one can consider applying the injection only in decoder layers (D1-D3) and after $5$ time steps. However, if the user wishes to maximize adaptation near seams and to maximize local appearance fidelity, we recommend using our full injection approach.
Appendix F Baseline Implementation Details
Below, we describe the implementation details of the baselines.
Interactive Digital Photomontage [1].   For their pixel-space graph cut, we used their “match edge” seam objective, where the pairwise term $E(p,q,L_{p},L_{q})$ is computed based on a Sobel filter on the RGB values. As shown in Figure 3, the edge strength computed from the filter may be sensitive to low-level changes in color, which could cause their graph cut to select seams that fall on undesired edges. Diffusion feature-based graph cut, on the other hand, selects seams that are more aligned with semantic features.
After their pixel-space graph cut, we followed their approach and used Poisson blending [64] as a post-process for smoothing composite regions together. Because generated images [64] tend to have wide color variations, gradient-domain blending often leads to changes in color, such as the marked regions in Figure 20. Our method is better at preserving local color and image appearances.
Blended Latent Diffusion [4].   BLD blends images by combining their noise at each diffusion step. For this baseline, we inject the noises of image $i=2...N$ into the noise of the base image $i=1$. As shown in Figure 20, BLD may change the appearance of the base image (a, c, e), the injected regions (d), and include artifacts (b, e, f).
@@ Appendix F Baseline Implementation Details
Blended Latent Diffusion + MultiDiffusion [4, 8].   We experimented with a modified version of Blended Latent Diffusion, where we fused the noises of different image regions with greater overlap, inspired by MultiDiffusion. Specifically, we dilate the mask of each region (in feature space) with a $3\times 3$ kernel and average the noise within the overlapped regions. As shown in Figure 21, this leads to artifacts near the seams (site of overlap) or changes in local appearances.
MasaCtrl with ControlNet [15].   We adapt their mask-guided framework by designating the base image as the target image and the other images within the stack as source images. We extend their framework to use multiple foreground masks (one for each image beyond the base image). As shown in Figure 20, this may change the appearance of the base image (a, c, e), selected regions (d, e, f), or show other blending artifacts (b).
CollageDiffusion [69].   We followed their released demo for generating baseline results. Their method allows users to tweak parameters per image layer and scene, such as added noise level and cross-attention modulation, and then outputs a composite. We created the input image layers by resizing the masks from feature-space graph-cut to image space and applying them to each image. Reducing the added noise to zero preserves all local regions but with little harmonization (similar to copy-paste). To minimize changes, we used a set of parameters that preserves the background region (base image) but allows the foreground injected regions to change. We set noise levels to $0.05$ for the base image and $0.4$ for the non-base images. We set noise blur to $30$ to smooth the seams. To ensure spatial fidelity of the image regions, we set the cross-attention modulation of non-base images to $0.5$, following their released examples. We also use their textual inversion feature and learn a special embedding for each image layer. If there are multiple image layers (regions) that correspond to the same word (e.g., multiple robot parts correspond to “robot” in the prompt “A robot from the future”), we choose the one that most closely matches the word description.
@@
Deep Image Blending [89].   We ran their released code to generate the baseline results. Their method takes in two images to composite (a source and target image) and a mask of the source image. To create the mask for each image, we resized the corresponding graph-cut mask from feature to image space. For composites of more than two input images, we iteratively ran Deep Image Blending on pairs of images to build toward the final composite. One input constraint is that their method takes square images as input, so we pre-process our image stack by resizing them to square images and resizing the output back to the original aspect ratios.
GP-GAN [85].   We ran their released code with their pre-trained model to generate the results. GP-GAN takes in two images to composite (a source and destination image), along with a mask of the source image to insert into the destination image. For composites of more than two images, we iteratively ran GP-GAN on pairs of images to build toward the final composite. We created the input masks by resizing the diffusion-space graph-cut results to the original resolution of each image.
@@
Appendix G User Survey Details
@@
Figure 12: User Survey (Blending). We show users the input images and selected regions to blend, and then ask them pick the best result from four images: one generated by our method, and the others generated by Interactive Digital Photomontage, Blended Latent Diffusion, and MasaCtrl+ControlNet.
@@
We launched two anonymous user surveys, in order to compare our method with our three most competitive baselines (Interactive Digital Photomontage [1], Blended Latent Diffusion [4], MasaCtrl+ControlNet [15]). Across $12$ scenes (Figures 13a-f, 15a-e, 16), we created results using our method and the three baselines. For each scene, we showed the four results side-by-side and asked users to select the best result. The order of the four images is randomized for each scene, and the order of scenes is also randomized in each survey.
@@
In the first survey, we asked “Which of the following images appear most realistic to you?” In the second survey, we showed the input images and asked users to pick the image that was the best at blending them. Specifically, we showed the full input images as well as the graph-cut regions (to be blended) on each input image. To help users focus on the individual image regions, we darken out other parts of the input image (Figure 12). Then, we asked “Which of the following images is BEST at blending all the selected image regions shown on top?”
@@
Users are asked to complete the first survey (realism) before the second survey (blending) to avoid bias. We received $27$ completed surveys for the realism one (a total of $324$ responses across all scenes) and $20$ completed surveys for the blending one (a total of $240$ responses across all scenes).
Appendix H Result Details
@@
We used various Stable Diffusion pre-trained models (v1.5) to generate results in this paper. Specifically, we used the canny edge model (Figures 13d-e, 15a,d,e, 16), the scribble model (Figures 13b,c,f, 15b, f), the depth map model (Figure 13a, g, 14b, 18), the Openpose model (Figures 14a, 15c), and the HED model (Figures 14c, 17, 19, 28).
@@
The input conditions (e.g., edge map, sketches) are manually created or derived from images released by ControlNet [90] or found on the web [48, 42, 84, 74, 20, 3, 54, 46, 52, 10, 59].
Appendix I Limitations
While we have shown our method’s versatility in various applications, we also observe several limitations. First, our current graph cut parameters are empirically chosen to encourage congruous regions, which penalizes seam circumference. While this works well for many cases, if the target object has a curvy outline, it may require additional user strokes to obtain a finer boundary (Figure 27). Since graph cut is solved in near real-time (${\sim}1$s), users can quickly check the graph-cut result and iterate as needed.
Second, our method assumes some spatial consistency among images in the stack. If the images differ significantly in scene structure, it will rely more on the user to select proper regions to form a valid scene (Figure 28a). Alternatively, users can increase spatial consistency by adding more spatial structure to the input control (Figure 28b). Future work can investigate ways to relax spatial consistency constraints and automatically account for dramatic scene structure changes during segmentation and blending.
Appendix J SDXL Results
Our initial tests show that our method works on SDXL as well. Please see Figure 29 for example results.
For SDXL, we kept all hyperparameters the same, except for the graph cut parameter $\sigma$, which we increased to $25$. We found that using the same $\sigma$ as in SD1.5 results in more cuts and less coherent segmentations in SDXL. This may be due to SDXL’s larger feature map size and differences in how feature variations correspond to semantic boundaries.
Figure 13: Results: Appearance Mixing. Examples of using Generative Photomontage for creative design. (a, b) Users can combine different architectural elements to form new architectural designs. (c, d, e) The user combines different vibrant colors of snakes or birds to explore new looks. (f, g) The user combines different parts of a futuristic robot and shoes to form their favorite look.
Figure 14: Results: Appearance Mixing (cont.). More examples of using our method for novel artistic and graphic designs. Users can (a) mix and combine iconic features of superhero costumes to form a new one, (b) mix different visual features when exploring artistic designs of a rocket, and (c) select and combine desired elements when designing a poster.
Figure 15: Results: Shape and Artifacts Correction. Our method can fix incorrect shapes and artifacts from ControlNet’s outputs, which often occur for uncommon input shapes. For example, (a) users can remove the extra rock at the Apple bite with a patch of grass from the second image, (b) correct the shape and contour of the first waffle by selecting the background region of the second image, and (f) refine the hand-shaped island in the second image with patches from the first image. Users can also (c) correct the dancer’s pose, (d) remove the extra leg of the dog, and (e) and replace the first dog with artifacts with the second dog.